OpenStudy (johnnydicamillo):

Given that f(x) =x^8h(x) h(-1) = 2 and h (-1) = 5 calculate f (-1)?

OpenStudy (johnnydicamillo):

so given that \[f(x) = x^8h(x), h(-1) =2, h(-1) =5 \] what is \[f'(-1)\]

jimthompson5910 (jim_thompson5910):

I'm assuming one of those h(-1) is h ' (-1)

OpenStudy (johnnydicamillo):

yes sorry

jimthompson5910 (jim_thompson5910):

which one

OpenStudy (johnnydicamillo):

the second one is \[h'(-1) = 5\]

jimthompson5910 (jim_thompson5910):

you first use the product rule, then you plug in x = -1 \[\Large f(x) = x^8*h(x)\] \[\Large f'(x) = 8x^7*h(x) + x^8*h'(x)\] \[\Large f'(-1) = 8(-1)^7*h(-1) + (-1)^8*h'(-1)\] \[\Large f'(-1) = ???\]

OpenStudy (johnnydicamillo):

-11?

jimthompson5910 (jim_thompson5910):

getting the same